OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how to use the distributive property to simplify the expression below. 2(x + 3) Part 2: Write the simplified expression.

OpenStudy (anonymous):

whoever gets it first gets a medal

OpenStudy (anonymous):

\[2\times \left( x+3 \right) = 2\times x + 2\times3 = 2x +6\]